340 tickets divide by 3
A has 80 more than B
B has 40 more than C
How many tickets do each have?

Answer by rfer(16322)

c=x
b=x+40
a=x+40+80
-----------
x+x+40+x+40+80=340
3x+160=340
3x=340-160
x=180/3
x=60 C
x+40=100 B
x+120=180 A